phone calls arrive at the rate of 30 per hour at the desk for regional airways 1) compute t he probability of recieving 2 calls in a 5 minute interval of time 2) compute the probality of recieving exactly 8 calls in 15 minutes
Answer #1

phone calls arrive at the rate of 30 per hour

part 1)

X :- Number of calls arrived.

X ~ P(λ 5/60 * 30 = 2.5)

P(X = 2)-(e- 2.5 (2.5)2)/2! = 0.2565

Part 2)

X ~ P(λ-15/60 * 30-75)

PIX = 8) = (e-1.5(7.5))/8! = 0.1373